# SPECIAL RELATIVITY

### v11

@@ 4.10 Spatiotemporal Transformation of the Transmitted Light Beam
Suppose the transmitted light pulse propagates in the $x-z$ plane. Our focus is on the space-time intensity distribution within this plane. Spatiotemporal coupling naturally emerges in the transmitted radiation behind the screen due to the angular-frequency dispersion introduced during the transmission process. In most cases, the emitted light beam can be accurately represented as the product of independent spatial and temporal factors. However, this assumption breaks down when the light must pass through an aperture in a moving opaque screen, as the transmission process disrupts the simple separation of space and time dependencies.
@@
We begin by expressing the field of an emitted pulse as $E=b_{i}(x)\exp[i\omega_{i}(z/c-t)]$. The initial amplitude distribution, $b_{i}(x)$, in front of the moving screen is the optical replica of the emitter’s aperture. The electric field of the transmitted pulse can be represented in the reciprocal domain as $\bar{E}(\Delta k_{x},\Delta\omega)=\bar{E}(K_{\perp},K_{\perp}d\omega/dk_{x})$, which corresponds to the Fourier component of a beam with angular frequency dispersion, where $d\omega/dk_{x}=v$. Taking the inverse Fourier transform, we obtain $E=b(x-vt)\exp[i\omega_{i}(z/c-t)]$. This represents the field immediately behind the moving screen. The beam profile, $b(x-vt)$, is the optical replica of the moving aperture. Now, consider a screen positioned at rest at a distance $l$ behind the moving aperture. For simplicity, we assume that the Fresnel number is large, $N_{F}=D_{h}^{2}\omega/(cl)\gg 1$, allowing us to neglect diffraction effects. Here, $D_{h}$ is the characteristic aperture size. From this, we conclude that the light spot on the observer’s screen moves with the same velocity $v$ as the moving screen. Additionally, we observe that a particular quantity remains unchanged: the combination $x-vt$ is the same for both the spot and the aperture. What does this imply? A careful examination reveals that events occur simultaneously at two separate locations along the $z$-axis  [^16] Spatiotemporal coupling is typically discussed in the literature in the context of ultrashort laser pulse propagation through a grating monochromator. In this scenario, in addition to considering phase fronts, one must also account for planes of constant intensity—known as pulse fronts. In a grating monochromator, the different spectral components of the outgoing pulse propagate in different directions. The electric field of a pulse with angular dispersion can be expressed in the Fourier domain $[\Delta k_{x},\Delta\omega]$ as $\bar{E}(\Delta k_{x}-p\Delta\omega,\Delta\omega)$. Its inverse Fourier transform to the space-time domain $[x,t]$ results in $E(x,t+px)$, representing a pulse with a pulse-front tilt. In our case of interest, we consider light diffracted by a tangentially moving set of gratings, where the Doppler effect induces frequency dispersion given by $d\omega/dk_{x}=v$. Consequently, the spatiotemporal coupling due to light transmission through a hole in a moving screen differs significantly from the conventional pulse-front tilt distortion..
@@ 4.11 Applicability of Ray Optics
Some experts argue that the applicability of the corpuscular model in the theory of light should be reinterpreted as the applicability of ray optics.  [^17] To quote Brillouin [25]: ”Fig. 5 explains the situation assuming a simplified device consisting of a parallel plate moving with uniform velocity $v$ in the horizontal direction. Monochromatic light falls normally on the plate and generates an oblique ray.” This oblique effect is demonstrated in Fig. 10.. Let us examine what happens, according to ray optics, in our case of interest. When light rays fall normally on a moving screen, they generate an oblique ray beam, as illustrated in Fig. 10. In this scenario, treating light rays as tiny particles leads to a well-known effect.
@@ 4.13 A Point Source in an Inertial Frame of Reference
The diffraction of a source field is typically categorized into Fresnel (near-zone) and Fraunhofer (far-zone) diffraction. In Fraunhofer diffraction, the wave phase is assumed to vary linearly across the detector aperture, which occurs, for example, when a plane wave impinges on the aperture at an angle to the optical axis. In contrast, Fresnel diffraction replaces this linear phase variation with a quadratic one. In the far zone, both types of sources effectively produce a plane wave in front of the pupil detection system. This implies that, in the Fraunhofer zone, the instrument always exerts a physical influence on the measurement of light aberration.
@@
One of the key conclusions of the discussion above (see Fig. 11) is that the aberration of a moving point source is absent in an inertial frame (Fig. 12). This leads to several remarkable effects, stemming from the fact that the transmitted light beam does not carry information about the motion of the point source. In fact, this principle is central to the binary star paradox discussed in Chapter 6.
To further explore the aberration of light in an inertial frame, it is useful to consider a slightly modified source—a ”plane wave” emitter. We adopt the model described in Section 4.2 because it is both relatively simple and sufficiently general to serve as a prototype for understanding aberration phenomena. While no one has performed all the ”thought experiments” exactly as described here, the outcomes can be confidently predicted based on the laws of special relativity, which are themselves grounded in experimental evidence.
@@ 4.15 Differences between the Light and Sound Aberration
Having completed our discussion on the aberration of light in an initial inertial frame, we now turn to optical phenomena in an accelerated frame. However, before proceeding, we will first explore optical-acoustic analogies. Sound is a periodic motion of air caused by vibrations of a source. The effect of sound aberration is best understood using a model of a single plane-wave emitter. Specifically, we consider a two-dimensional array of identical, coherent elementary sources uniformly distributed on a given $(x-y)$ plane $P$, all emitting waves simultaneously. We analyze a scenario in which a finite-aperture mirror moves tangentially to its surface. For simplicity, we assume the mirror’s size is small compared to the plane-wave emitter. An observer, at rest in the atmosphere frame, describes the emitted sound beam using the diagonal wave equation: $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$, where $v_{s}$ is the speed of sound.
@@ 5.1 Absolute Time Coordinatization in Accelerated Systems
We analyze the problem of light aberration in accelerated systems within the framework of special relativity. Specifically, we demonstrate that explaining optical effects in a rotating frame of reference does not require modifying special relativity or invoking general relativity. Instead, a strict adherence to special relativity suffices. To express electrodynamics laws in an accelerated frame, one must take an additional step by defining the metric of the non-inertial frame. Importantly, the metric tensor must remain a continuous function, leading to the concept of smoothly tailoring the metric tensor.
This problem in special relativity can be effectively addressed using an approach based on absolute time coordinatization. At the end of this section, we outline the mathematical extension of special relativity to accelerating systems. However, in some cases, direct physical arguments can provide quicker insights while remaining consistent with rigorous mathematical derivations.
Suppose the initial inertial system $S$ is at rest with respect to the fixed stars, and a system $S_{n}$ (along with an observer and their measuring instruments) is accelerated from rest in $S$ to a velocity $\vec{v}$ along the $x$-axis. In accelerated systems, the concept of maintaining absolute simultaneity remains logically consistent with the smooth tailoring of the metric tensor. The synchronization method involves preserving, without modification, the same set of uniformly synchronized clocks used when $S_{n}$ was initially at rest. It is well known that during acceleration (relative to the fixed stars), Einstein’s clock synchronization procedure cannot be applied. Consequently, by the time $S_{n}$ transitions to uniform motion, the metric in the accelerated reference frame will have acquired a non-diagonal form.
We begin by considering the metric as the true measure of space-time interval for an accelerated observer $S_{n}$ with coordinates $(t_{n},x_{n},y_{n},z_{n})$. To transform to the coordinates of an inertial observer $S$ moving with velocity $-v$ relative to $S_{n}$, we apply the inverse Galilean transformation $x=x_{n}+vt$, while leaving the time coordinate unchanged, $t_{n}=t$. Substituting this into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$, we obtain the transformed metric:
@@
This metric describes the electrodynamics of the accelerated light source with the viewpoint of the accelerated observer measurements as viewing this of the inertial observer. Inspecting Eq.(15) we can find the components of the metric tensor $g_{\mu\nu}$ in the coordinate system of $S_{n}$. We obtain $g_{00}=1-v^{2}/c^{2}$, $g_{01}=-v/c$, $g_{11}=-1$. Note that the metric in Eq. (15) is not diagonal, since, $g_{01}\neq 0$, and this implies that time is not orthogonal to space. This result represents the Langevin metric, derived by matching the metric tensors of the accelerated frame and the initial inertial frame. Considering the Langevin metric, we conclude that the length of a physical rod $dx$ and the proper time interval $dt$ in the inertial frame will coincide with the coordinate distance and coordinate time in the accelerated frame, i.e., $dx_{n}=dx$, $dt_{n}=dt$. (For further details, see Chapter 12).
@@
Applying the inverse Galilean transformation, we have $dt_{n}=dt$, $dx=dx_{n}+atdt/\sqrt{1+a^{2}t^{2}/c^{2}}$. With this transformation, the metric in the uniformly accelerating observer’s frame takes the form:
@@ 5.2 Asymmetry Between Inertial and Accelerated Frames
The velocity of light emitted by a source at rest in the coordinate system $(t,x,y,z)$ of the inertial frame $S$ is $c$. The Minkowski metric, given by Eq. (1), implies a symmetry in the one-way speed of light in this frame. However, in the coordinate system $(t_{n},x_{n},y_{n},z_{n})$ of an accelerated frame $S_{n}$, the speed of light emitted by a source at rest in $S_{n}$ differ from $c$ This is because $(t_{n},x_{n},y_{n},z_{n})$ is related to $(t,x,y,z)$ via a Galilean transformation. This is readily verified if one recalls that the velocity of light in the reference system $S$ is equal to $c$. If $ds$ is the infinitesimal displacement along the world line of a ray of light, then $ds^{2}=0$ and we obtain $c^{2}=(dx/dt)^{2}$. In the accelerated reference system, since $x_{n}=x-vt$ and $t=t_{n}$, this expression takes the form $c^{2}=(dx_{n}/dt_{n}+v)^{2}$, as seen by setting $ds^{2}=0$ in Eq. (15). Consequently, in the accelerated coordinate system $(ct_{n},x_{n})$, the velocity of light parallel to the x-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction, as expected.
@@
It is useful to begin with an overview of some key results. Consider an observer in the accelerated frame $S_{n}$ performing an aberration measurement. To predict to the outcome, the observer must use the non-diagonal metric given in Eq. (15). Due to the inherent asymmetry between inertial and accelerated frames, an intriguing prediction arises in the theory of light aberration. Specifically, if an opaque screen with a hole is initially at rest relative to the fixed stars and then set into motion, the apparent angular position of a ”plane-wave” emitter observed through the aperture in the accelerated frame will shift by an angle $-v/c$. This effect is illustrated in Fig. 14. Notably, the source of this asymmetry is not the relative acceleration between the two frames but rather the difference in their individual acceleration histories with respect to the fixed stars.
@@ 5.3 Electrodynamic Explanation of Light Aberration
where coordinates and time are transformed according to a Galilean transformation. Acceleration affects the propagation of light. The coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c-v$ in the positive direction, and $dx/dt=-c-v$ in the negative direction.
@@
Acceleration influences the field equations, transforming the wave equation into Eq. (17). Consider a transmitted plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$. With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ equation Eq.(17) becomes: $(1-v^{2}/c^{2})k_{x}^{2}-2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$. The wavenumber vector of the transmitted plane wave is fixed by initial conditions. Suppose the wavefront of the emitted light beam is initially perpendicular to the vertical direction before acceleration. Then, we have $k_{z}=-\omega_{i}/c$, $k_{x}=K_{\perp}$, where $K_{\perp}$ is the wavenumber accosiated with space-modulated transmittance.
[…]
@@
Now, let us address an apparent paradox. We derived the electrodynamics equations in an accelerated frame using the Galilean transformation of Maxwell’s equations. Intuitively, one might expect everything to be at rest in the accelerated frame. However, special relativity dictates the presence of a time derivative, $\partial/\partial t_{n}=v\partial/\partial x_{n}$, since $\partial/\partial x_{n}$ is nonzero. This situation differs fundamentally from electrodynamics in an inertial frame. We will explore this topic further in Chapter 9.

### v12

@@ 4.10 Spatiotemporal Transformation of the Transmitted Light Beam
Suppose the transmitted light pulse propagates in the xzx-z plane. Our focus is on the space-time intensity distribution within this plane. Spatiotemporal coupling naturally emerges in the transmitted radiation behind the screen due to the angular-frequency dispersion introduced during the transmission process. In most cases, the emitted light beam can be accurately represented as the product of independent spatial and temporal factors. However, this assumption breaks down when the light must pass through an aperture in a moving opaque screen, as the transmission process disrupts the simple separation of space and time dependencies.
@@
We begin by expressing the field of an emitted pulse as $E=b_{i}(x)\exp[i\omega_{i}(z/c-t)]$. The initial amplitude distribution, $b_{i}(x)$, in front of the moving screen is the optical replica of the emitter’s aperture. The electric field of the transmitted pulse can be represented in the reciprocal domain as $\bar{E}(\Delta k_{x},\Delta\omega)=\bar{E}(K_{\perp},K_{\perp}d\omega/dk_{x})$, which corresponds to the Fourier component of a beam with angular frequency dispersion, where $d\omega/dk_{x}=v$. Taking the inverse Fourier transform, we obtain $E=b(x-vt)\exp[i\omega_{i}(z/c-t)]$. This represents the field immediately behind the moving screen. The beam profile, b(xvt)b(x-vt), is the optical replica of the moving aperture. Now, consider an observer’s screen positioned at rest at a distance $l$ behind the moving aperture. For simplicity, we assume that the Fresnel number is large, $N_{F}=D_{h}^{2}\omega/(cl)\gg 1$, allowing us to neglect diffraction effects. Here, $D_{h}$ is the characteristic aperture size. From this, we conclude that the light spot on the observer’s screen moves with the same velocity $v$ as the moving screen. Additionally, we observe that a particular quantity remains unchanged: the combination xvtx-vt is the same for both the spot and the aperture. What does this imply? A careful examination reveals that events occur simultaneously at two separate locations along the $z$-axis.  [^16] Spatiotemporal coupling is typically discussed in the literature in the context of ultrashort laser pulse propagation through a grating monochromator. In this scenario, in addition to considering phase fronts, one must also account for planes of constant intensity—known as pulse fronts. In a grating monochromator, the different spectral components of the outgoing pulse propagate in different directions. The electric field of a pulse with angular dispersion can be expressed in the Fourier domain $\Delta k_{x},\Delta\omega]$ as $\bar{E}(\Delta k_{x}-p\Delta\omega,\Delta\omega)$. Its inverse Fourier transform to the space-time domain $[x,t]$ results in $E(x,t+px)$, representing a pulse with a pulse-front tilt. In our case of interest, we consider light diffracted by a tangentially moving set of gratings, where the Doppler effect induces frequency dispersion given by $d\omega/dk_{x}=v$. Consequently, the spatiotemporal coupling due to light transmission through a hole in a moving screen differs significantly from the conventional pulse-front tilt distortion.
@@ 4.11 Applicability of Ray Optics
Some experts argue that the applicability of the corpuscular model in the theory of light should be reinterpreted as the applicability of ray optics.  [^17] To quote Brillouin [25]: ”Fig. 5 explains the situation assuming a simplified device consisting of a parallel plate moving with uniform velocity $v$ in the horizontal direction. Monochromatic light falls normally on the plate and generates an oblique ray.” This oblique effect is demonstrated in Fig. 10. Let us examine what happens, according to ray optics, in our case of interest. When light rays fall normally on a moving screen, they generate an oblique ray beam, as illustrated in Fig. 10. In this scenario, treating light rays as tiny particles leads to a well-known effect.
@@ 4.13 A Point Source in an Inertial Frame of Reference
The diffraction of a source field is typically categorized into Fresnel (near-zone) and Fraunhofer (far-zone) diffraction. In Fraunhofer diffraction, the wave phase is assumed to vary linearly across the detector aperture, which occurs, for example, when a plane wave impinges on the aperture at an angle to the optical axis. In contrast, Fresnel diffraction replaces this linear phase variation with a quadratic one. In the far zone, both types of sources effectively produce a plane wave in front of the pupil detection system.
@@
Let us now consider the case of a moving point source. Specifically, we examine a setup where the screen is at rest, while the point source moves tangentially. A key takeaway from the earlier discussion is that no aberration is observed for a large-aperture plane wave emitter moving tangentially in an inertial frame (see Fig. 11).
Remarkably, the same conclusion holds for a point source. This stems from the linearity of Maxwell’s equations: the governing laws of electrodynamics are linear differential equations. As discussed previously, the most effective approach to analyzing a moving emitter is via the Fourier transform method. A point source can be decomposed into a superposition of spatial Fourier components—each corresponding to a radiated plane wave.
The current density associated with these components can be modeled as $j_{pol}=j_{0}\sin{({\vec{K}}_{\perp}\cdot\vec{r})}$. Each elementary dipole source radiates a plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$, where the wavevector $\vec{k}$ is determined by the initial spatial modulation. Near normal incidence, we set $k_{x}=K_{\perp}$, where $K_{\perp}$ is the transverse spatial frequency corresponding to the modulated dipole density. For small angles, the incidence angle is approximately given by $\theta\simeq K_{\perp}/k_{z}\simeq K_{\perp}c/\omega$.
[…]
This leads to a remarkable conclusion: the transmitted beam carries no information about the source’s tangential motion. This principle lies at the heart of the binary star paradox discussed in Chapter 6.
@@ 4.15 Differences between the Light and Sound Aberration
Having completed our discussion on the aberration of light in an initial inertial frame, we now turn to optical phenomena in an accelerated frame. However, before proceeding, we will first explore optical-acoustic analogies. Sound is a periodic motion of air caused by vibrations of a source. The effect of sound aberration is best understood using a model of a single plane-wave emitter. Specifically, we consider a two-dimensional array of identical, coherent elementary sources uniformly distributed on a given (xy)(x-y) plane $P$, all emitting waves simultaneously. We analyze a scenario in which a finite-aperture mirror moves tangentially to its surface. For simplicity, we assume the mirror’s size is small compared to the plane-wave emitter. An observer, at rest in the atmosphere frame, describes the emitted sound beam using the diagonal wave equation: $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$, where $v_{s}$ is the speed of sound.
@@ 5.1 Absolute Time Coordinatization in Accelerated Systems
We investigate the phenomenon of light aberration in accelerated systems within the framework of special relativity. In particular, we demonstrate that explaining optical effects in rotating frames does not require modifying special relativity or invoking general relativity. A rigorous application of special relativity is sufficient. However, to express electrodynamics in a non-inertial frame, one must take the additional step of defining the metric of that frame. Crucially, the metric tensor must remain a continuous function, leading naturally to the concept of smoothly tailoring the metric.
This issue in special relativity can be effectively addressed using an approach based on absolute time coordinatization. Toward the end of this section, we outline how special relativity may be mathematically extended to describe accelerating systems. In some situations, however, direct physical reasoning can yield rapid insights while remaining consistent with formal derivations.
Consider an inertial frame $S$, assumed to be at rest relative to the fixed stars. A second frame, $S_{n}$, along with an observer and measuring instruments, is accelerated from rest in $S$ to a velocity $v$ along the $x$-axis. In such accelerated systems, the notion of absolute simultaneity remains logically consistent with the smooth tailoring of the metric tensor. Synchronization in this context involves preserving the same uniformly synchronized clocks that were used when $S_{n}$ was initially at rest. It is well known that during acceleration relative to the fixed stars, Einstein’s synchronization procedure cannot be directly applied. As a result, by the time $S_{n}$ reaches uniform motion, the metric in its reference frame will generally acquire a non-diagonal form.
We begin by treating the metric as the true measure of spacetime intervals for an accelerated observer in frame $S_{n}$, with coordinates $(t_{n},x_{n},y_{n},z_{n})$. To express this in the coordinates of an inertial observer in $S$, moving at velocity v-v relative to $S_{n}$, we apply the inverse Galilean transformation $x=x_{n}+vt$, keeping the time coordinate unchanged: $t_{n}=t$. Substituting into the Minkowski metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$, we obtain the transformed (Langevin) metric:
@@
It is crucial to note that Langevin metric, in fact, reflects measurements performed by an inertial observer using the Minkowski metric, subsequently re-expressed via the transformation $x\to x_{n}+vt$, $t\to t_{n}$. This metric describes the electrodynamics of the accelerated light source with the viewpoint of the accelerated observer measurements as viewing this of the inertial observer.
Inspecting Eq.(15) we can find the components of the metric tensor $g_{\mu\nu}$ in the coordinate system of $S_{n}$. We obtain $g_{00}=1-v^{2}/c^{2}$, $g_{01}=-v/c$, $g_{11}=-1$. Note that the metric in Eq. (15) is not diagonal, since, $g_{01}\neq 0$, and this implies that time is not orthogonal to space. This result represents the Langevin metric, derived by matching the metric tensors of the accelerated frame and the initial inertial frame.
To interpret Eq. (15) physically, recall that a new frame of reference can always be introduced via a passive coordinate transformation—a mere relabeling of coordinates used to describe physical events. For example, an observer in the inertial frame can introduce a comoving coordinate system to analyze radiation from a moving source. In this system, fields are expressed as functions of $(t_{n},x_{n},y_{n},z_{n})$, which relate to the original coordinates $(t,x,y,z)$ through a passive Galilean transformation.
From the structure of the Langevin metric, it is evident that the electromagnetic field measured at position $x_{n}$ and time $t_{n}$ by a device at rest in the accelerated frame will match the field measured by an identical device at rest in the inertial frame at position $x=x_{n}+vt$ and time $t=t_{n}$. Substituting this coordinate relationship into the Minkowski metric of the inertial observer leads directly to Eq. (15).
In describing physical phenomena in an accelerated frame, one must distinguish between coordinate quantities and physical quantities. The Langevin metric allows us to determine how these relate. For instance, the length of a physical rod $dx$ and the proper time interval $dt$ in the inertial frame coincide with the coordinate length and time in the accelerated frame: $dx_{n}=dx$, $dt_{n}=dt$. For a more detailed discussion on how to relate coordinate and physical quantities in non-inertial frames, refer to Chapter 12.
@@
Applying the inverse Galilean transformation, we have $dt=dt_{n}$, $dx=dx_{n}+at_{n}dt_{n}/\sqrt{1+a^{2}t_{n}^{2}/c^{2}}$. With this transformation, the metric in the uniformly accelerating observer’s frame takes the form:
@@ 5.2 Asymmetry Between Inertial and Accelerated Frames
The velocity of light emitted by a source at rest in the coordinate system $(t,x,y,z)$ of the inertial frame $S$ is $c$. The Minkowski metric, given by Eq. (1), implies a symmetry in the one-way speed of light in this frame. However, in the coordinate system $(t_{n},x_{n},y_{n},z_{n})$ of an accelerated frame $S_{n}$, the speed of light emitted by a source at rest in $S_{n}$ differ from $c$ This is because $(t_{n},x_{n},y_{n},z_{n})$ is related to $(t,x,y,z)$ via a Galilean transformation. This is readily verified if one recalls that the velocity of light in the reference system $S$ is equal to $c$. If $ds$ is the infinitesimal displacement along the world line of a ray of light, then $ds^{2}=0$ and we obtain $c^{2}=(dx/dt)^{2}$. In the accelerated reference system, since $x_{n}=x-vt$ and $t_{n}=t$, this expression takes the form $c^{2}=(dx_{n}/dt_{n}+v)^{2}$, as seen by setting $ds^{2}=0$ in Eq. (15). Consequently, in the accelerated coordinate system $(ct_{n},x_{n})$, the velocity of light parallel to the x-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction, as expected.
@@
It is useful to begin with an overview of some key results. Consider an observer in the accelerated frame $S_{n}$ performing an aberration measurement. To predict to the outcome, the observer must use the non-diagonal metric given in Eq. (15). Due to the inherent asymmetry between inertial and accelerated frames, an intriguing prediction arises in the theory of light aberration. Specifically, if an opaque screen with a hole is initially at rest relative to the fixed stars and then set into motion, the apparent angular position of a ”plane-wave” emitter observed through the aperture in the accelerated frame will shift by an angle v/c-v/c. This effect is illustrated in Fig. 13. Notably, the source of this asymmetry is not the relative acceleration between the two frames but rather the difference in their individual acceleration histories with respect to the fixed stars.
@@ 5.3 Electrodynamic Explanation of Light Aberration
where coordinates and time are transformed according to a Galilean transformation. Acceleration affects the propagation of light. The coordinate velocity of light parallel to the $x_{n}$-axis is given by $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction.
@@
Acceleration influences the field equations, transforming the wave equation into Eq. (17). Consider a transmitted plane wave of the form $\exp(i\vec{k}_{n}\cdot\vec{r}_{n}-i\omega_{n}t_{n})$. With a plane wave $\exp(i\vec{k}_{n}\cdot\vec{r}_{n}-i\omega_{n}t_{n})$ with the wavenumber vector $\vec{k}_{n}$ and the frequency $\omega_{n}$ equation Eq.(17) becomes: $(1-v^{2}/c^{2})(k_{n})_{x}^{2}-2v(k_{n})_{x}\omega_{n}/c^{2}+(k_{n})_{z}^{2}-\omega_{n}^{2}/c^{2}=0$.
The wavevector $\vec{k}_{n}$ is determined by the initial conditions before the acceleration. Transforming to the accelerated frame using Galilean transformation $x=x_{n}+vt$, $t=t_{n}$ we find that the frequency shifts according to: $\omega_{n}=\omega_{i}-K_{\perp}v$, while the components of wavevector remain unchanged. Substituting these into dispersion relation confirms that it remains satisfied, as expected.
@@
Now, let us address an apparent paradox. We derived the electrodynamics equations in an accelerated frame using the Galilean transformation of Maxwell’s equations. From the Galilean transformation $x_{n}=x-vt,~t_{n}=t$, after partial differentiation, one obtains $\partial/{\partial t}=\partial/{\partial t_{n}}-v\partial/{\partial x_{n}}$, $\partial/{\partial x}=\partial/{\partial x_{n}}$. Intuitively, one might expect everything to be at rest in the accelerated frame. However, special relativity dictates the presence of a time derivative, $\partial/\partial t_{n}=v\partial/\partial x_{n}$, since $\partial/\partial x_{n}$ is nonzero. This situation differs fundamentally from electrodynamics in an inertial frame. We will explore this topic further in Chapter 9.